Problem 13-12 Prepare a Statement of Cash Flows [LO13-1, LO13-2]

A comparative balance sheet and an income statement for Burgess Company are given below:

Burgess Company
Comparative Balance Sheet
(dollars in millions)
Ending Balance Beginning Balance
Assets
Current assets:
Cash and cash equivalents $ 48 $ 99
Accounts receivable 730 669
Inventory 695 646
Total current assets 1,473 1,414
Property, plant, and equipment 1,595 1,565
Less accumulated depreciation 824 678
Net property, plant, and equipment 771 887
Total assets $ 2,244 $ 2,301
Liabilities and Stockholders' Equity
Current liabilities:
Accounts payable $ 278 $ 169
Accrued liabilities 189 161
Income taxes payable 95 81
Total current liabilities 562 411
Bonds payable 460 690
Total liabilities 1,022 1,101
Stockholders' equity:
Common stock 190 190
Retained earnings 1,032 1,010
Total stockholders' equity 1,222 1,200
Total liabilities and stockholders' equity $ 2,244 $ 2,301
Burgess Company
Income Statement
(dollars in millions)
Sales $ 3,970
Cost of goods sold 2,720
Gross margin 1,250
Selling and administrative expenses 898
Net operating income 352
Nonoperating items:
Gain on sale of equipment 2
Income before taxes 354
Income taxes 130
Net income $ 224

Burgess also provided the following information:

  1. The company sold equipment that had an original cost of $30 million and accumulated depreciation of $16 million. The cash proceeds from the sale were $16 million. The gain on the sale was $2 million.
  2. The company did not issue any new bonds during the year.

  3. The company paid a cash dividend during the year.

  4. The company did not complete any common stock transactions during the year.

Required:

Using the indirect method, prepare a statement of cash flows for the year. (Enter your answers in millions not in dollars. List any deduction in cash and cash outflows as negative amounts.)
Answer #1
Cash Flow Statement
Indirect Method
Cash flow from Operating Activities
Net Income $            224.00
Adjustments
Depreciation $          162.00 =824-(678-16)
Gain on sale of equipment $             -2.00
Change In current assets & Liabilities
Increase in Accounts Receivable $           -61.00 =669-730
Increase in inventories $           -49.00 =646-695
Increase in Accounts payable $          109.00 =278-169
Increase in Accrued Liabilities $            28.00 =189-161
Increase in Income Tax Payable $            14.00 =95-81
Total Adjustments $            201.00
Cash from operating activities $             425.00
Cash flow from Investing Activities
Cash paid for Purchase of Equipment $             -60.00 =1565-30-1595
Sale of Equipment $              16.00
Cash used in investing activities $             -44.00
Cash flow from Financing Activities
Repayment of Bonds Payable $          -230.00 =460-690
Dividend Paid $          -202.00 =-(1010+224-1032)
Cash from financing activities $           -432.00
Net Increase in cash $             -51.00
Opening Balance of Cash $               99.00
Closing Balance of Cash $               48.00